OpenStudy (jaylee):

what is 11 5/8 divided by 2

OpenStudy (anonymous):

5.8125

OpenStudy (dman):

5 13 16

OpenStudy (jaylee):

Thank You dman! Been way to long since doing fractions, thx for your help! Have a great day!

OpenStudy (dman):

ok you to

OpenStudy (texaschic101):

do you understand how the answer was reached ? 11 5/8 divided by 2 (turn the mixed number into an improper fraction) 93/8 divided by 2 ... 93/8 x 1/2 = 93/16 = 5 13/16 or 5.8125 (both answers are correct)